1. $39.00 per share is the current price for Foster Farms' stock. The dividend is projected to increase at a constant rate of 5.50% per year. The required rate of return on the stock, rs, is 9.00%. What is the stock's expected price 3 years from today?

Select the correct answer.

a. $46.19

b. $46.97

c. $46.58

d. $45.80

e. $45.41

2. Carby Hardware has an outstanding issue of perpetual preferred stock with an annual dividend of $5.50 per share. If the required return on this preferred stock is 6.5%, at what price should the preferred stock sell?

Select the correct answer.

a. $85.64
b. $83.60
c. $84.11
d. $85.13
e. $84.62

3. Orwell building supplies' last dividend was $1.75. Its dividend growth rate is expected to be constant at 25.00% for 2 years, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the current stock price?

Select the correct answer.

a. $41.98
b. $44.62
c. $42.64
d. $43.96
e. $43.3
Answer #1

1.P3=Current price*(1+Growth rate)^3

=39*(1.055)^3

=$45.8(Approx)

2.Current price=Annual dividend/required return

=5.5/0.065

=$84.62(Approx)

3.D1=(1.75*1.25)=2.1875

D2=(2.1875*1.25)=2.734375

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.734375*1.06)/(0.12-0.06)

=48.3072917

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.1875/1.12+2.734375/1.12^2+48.3072917/1.12^2

=$42.64(Approx)